A slip for tape casting contains, on a weight basis, 100 parts Al2O3, 10 parts nonvolatile organic, and 35 parts toluene. After drying, the adherent tap is 60% of the original thickness of the cast. Estimate the porosity of the tape after drying and after binder burnout. Assuming the nonvolatile organic has 1 g/cc density, all the organics are removed and no tape shrinkage occurs during binder burnout.
Answer #1

som Pжкоме #алла алды он: 100g Al2O3 log non volatie Organic 35 9 τοιυεου Aften drying an toluene is removed; Wdry= nog AftenVsoud = og/3959/4c = 278 cc porosity after drying = V void /vary -(63-300-27.822). 0.56 63-3 cc safe to anume 1. Uburn = V